# Adversarial Cooperative Rationalization: The Risk of Spurious Correlations in Even Clean Datasets

### v4

@@ 1 Introduction
Figure 2: A cherry-picked example of the generator-added spurious correlation. The underlined text is human-annotated rationale. The text in red is the rationale selected by RNP. Example 1: from a positive input $X^{1}$ with a label $1$, the generator selects a rationale $Z^{1}$ that includes the pattern “.”; and for a negative input $X^{0}$ with a label 00, the generator selects a rationale $Z^{0}$ that does not include“.”. And subsequently, the predictor considers the presence or absence of “.” as an indicative feature for positive classification.
Our research starts with a special empirical observation. We first observe that, even if we remove “maximizing the prediction accuracy” from the generator’s objective (thus it selects some random noise), the predictor can still be trained to get very high accuracy with these randomly selected spurious rationales (the orange line in Figure 3(a) of §$\S$4.1). This phenomenon then leads to a trust concern: whether the extracted rationale is really responsible for the label in the original dataset (i.e., although the extracted rationale is considered faithful to the model’s prediction by previous research, is it faithful to the model-agnostic dataset?). This problem is important because explanations should also be aligned with their social attribution (Jacovi & Goldberg, 2020, 2021).
@@ 3 Background of the rationalization task
Unless otherwise specified, uppercase letters represent random variables, while lowercase letters correspond to their values. For simplicity, we do not distinguish between vectors and scalars. We consider the classification task. We have a classification dataset $\mathcal{D}$, which can be seen as a collection of samples drawn from the true data distribution $P(X,Y)$. $X=X_{1:l}$ is the input text sequence of length $l$, and $Y$ represent the classes in the dataset (note that a discrete label can also seen as representing a distribution like [0,1]). By enumerating $X$, we can get $P(Y|X)$, which is the distribution that a normal non-interpretable classifier working on $\mathcal{D}$ needs to approximate. Rationalization consists of a generator $f_{g}(\cdot)$ (or ${g}$ for conciseness) and a predictor $f_{p}(\cdot)$, with $\theta_{g},\theta_{p}$ being their parameters.
@@
Figure 3: Experiments on the Beer-Aroma dataset (other datasets are in Appendix B.2): ``````` `full text"""": a predictor trained using the full texts. ``````` `random patterns"""": a predictor trained with randomly selected patterns. ``````` `r2f"""": feeding the random patterns to the predictor that was trained using the full texts.
@@
$\Omega(M)=\lambda_{1}\bigg{\lvert}\frac{||M||_{1}}{l}-s\bigg{\rvert}+\lambda_{2}\sum_{t=2}^{l}\big{|}M_{t}-M_{t-1}\big{|}.$ (4)
@@ 4 Motivation and method
Notation. For the sake of exposition, let us take the example of binary sentiment classification. We denote $X^{1}$ and $X^{0}$ as input texts with label $Y=1$ and $Y=0$, respectively. $Z$ and $Z_{A}$ represent the rationale candidates selected by the generator and the attacker, respectively. Note that they are not separate variables but a proxy for any variables within $X$. Sometimes we use $Z$ and the variable represented by $Z$ interchangeably. $T$ is a proxy for any variables within $\{T_{1},\cdots,T_{n}\}$ (defined in §$\S$3).
@@ 4.1 Cause of the spurious correlation
Example 1 in Figure 2 of §$\S$1 also provides an intuition for the above analysis.
@@
We provide more evidence of the existence of such spurious correlations in practical scenarios from another perspective by demonstrating the attack success rate in §$\S$5.2.
@@ 4.2 The proposed method
$Y_{A}$ represents the class to be attacked. We randomly select a class for each attack to create a balanced attack for each class. $[1/n,\cdots,1/n]$ represents the distribution of $P(Y)$ in the raw dataset. $\mathop{\min}_{\theta_{p}}H_{c}([1/n,\cdots,1/n],f_{p}(f_{a}(X)|f_{a}(X))$ means we rectify the sampled distribution of $P(Y|Z_{A},a)$ to $P(Y)$ and ask the predictor to learn that $Z_{A}$ is not correlated with $Y$. In binary classification, we have $Y_{A}=1-Y$ and $0.51/n=$ 0.5.
@@
During training, (9) and (8) are alternated. The practical implementation details with Pytorch are in Appendix A.3. The overall mechanism of the model is as follows: (8) inspects trivial patterns ($f_{a}(X)$) from $X$. The second term of (9) is the instruction that prevents the predictor from learning the trivial patterns by classifying them as random noise. A well instructed predictor is then able to give good feedback to the generator’s selection. And the first term of (9) is the normal RNP. The reason why the attacker constructed in this manner can detect trivial patterns will be elucidated in §$\S$4.3. We also use a toy example in Appendix A.2 to provide an intuitive understanding. At the end of §$\S$4.3, we also discuss how our method will work in the situation where the generator and the predictor cooperate correctly on real rationales rather than trivial patterns.
@@ 4.3 Underlying principles
Attack as inspection. Following the above settings for $Z=T$ and $I(Y;T)=0$ in §$\S$4.1, we will show how the trivial patterns learned by the predictor can be inspected through attack. Corresponding to (7), if the attack generator can be constructed in any way (i.e., has infinite expressiveness), then we can always find an attack generator $g_{a}$ which extracts $Z_{A}$ from $X$, such that
@@
Appendix C.1 shows the detailed derivation for the reason why we can find such a $g_{a}$. Equation (13) is the opposite of (7), and it means that under condition $g_{a}$, $T=t_{+}$ now becomes a negative class indicator, which is exactly the opposite situation under condition $g$. Here is the intuitive understanding of the attack. Corresponding to the punctuation pattern example mentioned in Figure 2 of §$\S$1. The generator $g$ selects $Z=``$."""" from $X^{1}$. And the predictor has learnt to predict ``."formulae-sequence``"``."` ` . " as positive. We can employ an attacker $g_{a}$ which selects $Z_{A}=``$."""" from $X^{0}$ (whose class label is negative) such that $Z_{A}$ can also be classified as positive. Similarly, the attacker can find $Z_{A}=``$,"""" from $X^{1}$ to be classified as negative. So, the overall objective of the attacker is to select those $Z_{A}$ that can be classified to the opposite class by the predictor.
@@
[…]
@@ 5.1 Settings
Baselines. We compare our A2I with the standard RNP and several recent representative methods: Inter_RAT (Yue et al., 2023) and CR (Zhang et al., 2023) represent recent causal methods, and FR (Liu et al., 2022) and NIR (Storek et al., 2023) represent recent methods designed to deal with degeneration. All of them have been discussed in §$\S$2.
@@ 6 Conclusion
The attachments are the new reference letters. Exploring XAI techniques can help to address many important problems existed in present deep learning models. For instance, XAI techniques can aid in detecting model discrimination (fairness) (Pradhan et al., 2022) or jailbreaks (Jiang et al., 2025), identifying backdoor attacks (security) (Yu et al., 2025a, b; Yang et al., 2025), controlling robots (Zhou et al., 2025), and revealing potential spurious correlations (robustness) (Yang et al., 2024), among others. Another direction is how to apply the findings of this direction into other modalities (Diao et al., 2024, 2025; Bi et al., 2025b, a).
@@ A.2 A toy example for a more intuitive understanding of the proposed method
Firstly, to inspect and identify the correlations, we introduce an attack generator $g_{a}$. Figure 7 shows an example of how the attacker works (formal analysis is in §$\S$4.3).
Example 2: the optimization objective of $g_{a}$ is to select an attack rationale $Z_{A}$ from input such that, when $Z_{A}$ is fed into the same predictor $p$, it yields a prediction label flipped from its original label. Continuing the previous example in Figure 2, the generator $g$ selects the “.” from a positive input $X^{1}$ with label $1$ as $Z$. Consequently, the predictor $p$ learns to treat the presence of “.” in $Z$ as an indicative feature for positive classification. On the other hand, the goal of $g_{a}$ is to select an attack rationale $Z_{A}$ from a negative input $X^{0}$ with a label 00 in such a way that, when $Z_{A}$ is fed to the same predictor $p$, the prediction result flips from its original label 00 to $1$. Achieving this objective is straightforward: $g_{a}$ simply needs to mimic $g$ by selecting “.” as $Z_{A}$. This suggests that if $g$ identifies $Z$ from $X^{1}$ as a trivial pattern also present in $X^{0}$, then $g_{a}$ can effortlessly select $Z_{A}=Z$ from $X^{0}$, leading to an easy flip of the prediction label of $Z_{A}$ to $1$ in predictor $p$. On the other hand, if $Z$ is a genuine positive rationale unique to $X^{1}$ and the predictor $p$ classifies it correctly, then $g_{a}$ would be unable to find a positive rationale from the negative input $X^{0}$. Therefore, it is difficult for the predictor $p$ to flip $Z_{A}$’s label from 00 to $1$. Thus, we can leverage the attack generator $g_{a}$ to assist in inspecting and identifying sampling bias. $g_{a}$ may easily find a $Z_{A}$ that flips its predicted label in predictor $p$ from its actual label, indicating the presence of semantically unrelated trivial patterns in $Z$.
@@ A.3 Implementation details of Equation (8) and (9)
Then, we get a copy of $Z_{A}$ with the pytorch function ``````` `torch.detach()"""":
@@
Note that this updating process will not influence the attacker, since we have used ``````` `torch.detach()"""" for $Z_{A}$.
@@ A.4 Datasets
The statistics of the datasets are in Table 5. $Pos$ and $Neg$ denote the number of positive and negative examples in each set. $S$ denotes the average percentage of tokens in human-annotated rationales to the whole texts. $avg\_len$ denotes the average length of a text sequence.
@@ C.1 Derivation of Equation (13)
The first property is a general property for conditional probability. If $10<P(Y=1)<$ 1, then for $\forall p$, if $10<p<$ 1, we can always find a variable $c$, such that $P(Y=1|c)=p$.
@@
If we can construct $G$ in an arbitrary way, and $10<P(Y=1|Z=t)<$ 1, then we have
@@ C.3 The minimum cross-entropy is equal to entropy
Figure 11: An example of llama’s output. Here “1” means that the class label $Y$ is positive. And the words after “||||” represent the rationale.

### v5

@@ 1 Introduction
Figure 2: A cherry-picked example of the generator-added spurious correlation. The underlined text is human-annotated rationale. The text in red is the rationale selected by RNP. Example 1: from a positive input $X^{1}$ with a label $1$, the generator selects a rationale $Z^{1}$ that includes the pattern “.”; and for a negative input $X^{0}$ with a label 0, the generator selects a rationale $Z^{0}$ that does not include“.”. And subsequently, the predictor considers the presence or absence of “.” as an indicative feature for positive classification.
Our research starts with a special empirical observation. We first observe that, even if we remove “maximizing the prediction accuracy” from the generator’s objective (thus it selects some random noise), the predictor can still be trained to get very high accuracy with these randomly selected spurious rationales (the orange line in Figure 3(a) of $\S$4.1). This phenomenon then leads to a trust concern: whether the extracted rationale is really responsible for the label in the original dataset (i.e., although the extracted rationale is considered faithful to the model’s prediction by previous research, is it faithful to the model-agnostic dataset?). This problem is important because explanations should also be aligned with their social attribution (Jacovi & Goldberg, 2020, 2021).
@@ 3 Background of the rationalization task
Unless otherwise specified, uppercase letters represent random variables, while lowercase letters correspond to their values. For simplicity, we do not distinguish between vectors and scalars. We consider the classification task. We have a classification dataset $\mathcal{D}$, which can be seen as a collection of samples drawn from the true data distribution $P(X,Y)$. $X=X_{1:l}$ is the input text sequence of length $l$, and $Y$ represent the classes in the dataset (note that a discrete label can also seen as representing a distribution like [0,1]). By enumerating $X$, we can get $P(Y|X)$, which is the distribution that a normal non-interpretable classifier working on $\mathcal{D}$ needs to approximate. Rationalization consists of a generator $f_{g}(\cdot)$ (or g{g} for conciseness) and a predictor $f_{p}(\cdot)$, with $\theta_{g},\theta_{p}$ being their parameters.
@@
Figure 3: Experiments on the Beer-Aroma dataset (other datasets are in Appendix B.2): ``full text": a predictor trained using the full texts. ``random patterns": a predictor trained with randomly selected patterns. ``r2f": feeding the random patterns to the predictor that was trained using the full texts.
@@
[…]
@@ 4 Motivation and method
Notation. For the sake of exposition, let us take the example of binary sentiment classification. We denote $X^{1}$ and $X^{0}$ as input texts with label $Y=1$ and $Y=0$, respectively. $Z$ and $Z_{A}$ represent the rationale candidates selected by the generator and the attacker, respectively. Note that they are not separate variables but a proxy for any variables within $X$. Sometimes we use $Z$ and the variable represented by $Z$ interchangeably. $T$ is a proxy for any variables within $\{T_{1},\cdots,T_{n}\}$ (defined in $\S$3).
@@ 4.1 Cause of the spurious correlation
Example 1 in Figure 2 of $\S$1 also provides an intuition for the above analysis.
@@
We provide more evidence of the existence of such spurious correlations in practical scenarios from another perspective by demonstrating the attack success rate in $\S$5.2.
@@ 4.2 The proposed method
$Y_{A}$ represents the class to be attacked. We randomly select a class for each attack to create a balanced attack for each class. $1/n,\cdots,1/n]$ represents the distribution of $P(Y)$ in the raw dataset. $\mathop{\min}_{\theta_{p}}H_{c}([1/n,\cdots,1/n],f_{p}(f_{a}(X)|f_{a}(X))$ means we rectify the sampled distribution of $P(Y|Z_{A},a)$ to $P(Y)$ and ask the predictor to learn that $Z_{A}$ is not correlated with $Y$. In binary classification, we have $Y_{A}=1-Y$ and $1/n=0.5$.
@@
During training, (9) and (8) are alternated. The practical implementation details with Pytorch are in Appendix A.3. The overall mechanism of the model is as follows: (8) inspects trivial patterns ($f_{a}(X)$) from $X$. The second term of (9) is the instruction that prevents the predictor from learning the trivial patterns by classifying them as random noise. A well instructed predictor is then able to give good feedback to the generator’s selection. And the first term of (9) is the normal RNP. The reason why the attacker constructed in this manner can detect trivial patterns will be elucidated in $\S$4.3. We also use a toy example in Appendix A.2 to provide an intuitive understanding. At the end of $\S$4.3, we also discuss how our method will work in the situation where the generator and the predictor cooperate correctly on real rationales rather than trivial patterns.
@@ 4.3 Underlying principles
Attack as inspection. Following the above settings for $Z=T$ and $I(Y;T)=0$ in $\S$4.1, we will show how the trivial patterns learned by the predictor can be inspected through attack. Corresponding to (7), if the attack generator can be constructed in any way (i.e., has infinite expressiveness), then we can always find an attack generator $g_{a}$ which extracts $Z_{A}$ from $X$, such that
@@
Appendix C.1 shows the detailed derivation for the reason why we can find such a $g_{a}$. Equation (13) is the opposite of (7), and it means that under condition $g_{a}$, $T=t_{+}$ now becomes a negative class indicator, which is exactly the opposite situation under condition $g$. Here is the intuitive understanding of the attack. Corresponding to the punctuation pattern example mentioned in Figure 2 of $\S$1. The generator $g$ selects Z=Z=``." from $X^{1}$. And the predictor has learnt to predict .``." as positive. We can employ an attacker $g_{a}$ which selects $Z_{A}=$``." from $X^{0}$ (whose class label is negative) such that $Z_{A}$ can also be classified as positive. Similarly, the attacker can find $Z_{A}=$``," from $X^{1}$ to be classified as negative. So, the overall objective of the attacker is to select those $Z_{A}$ that can be classified to the opposite class by the predictor.
@@
That is to say, although we cannot promise the independence between $Z_{A}$ and $Y$ under the generator’s conditional sampling, we can make $\upmodelsY^Z_{A}\upmodels\hat${Y} through the predictor’s prediction.
@@ 5.1 Settings
Baselines. We compare our A2I with the standard RNP and several recent representative methods: Inter_RAT (Yue et al., 2023) and CR (Zhang et al., 2023) represent recent causal methods, and FR (Liu et al., 2022) and NIR (Storek et al., 2023) represent recent methods designed to deal with degeneration. All of them have been discussed in $\S$2.
@@ 6 Conclusion
Exploring XAI techniques can help to address many important problems existed in present deep learning models. For instance, XAI techniques can aid in detecting model discrimination (fairness) (Pradhan et al., 2022) or jailbreaks (Jiang et al., 2025), identifying backdoor attacks (security) (Yu et al., 2025a, b; Yang et al., 2025), controlling robots (Zhou et al., 2025), and revealing potential spurious correlations (robustness) (Yang et al., 2024), among others. Another direction is how to apply the findings of this direction into other modalities (Diao et al., 2024, 2025; Bi et al., 2025b, a).
@@ A.2 A toy example for a more intuitive understanding of the proposed method
Firstly, to inspect and identify the correlations, we introduce an attack generator $g_{a}$. Figure 7 shows an example of how the attacker works (formal analysis is in $\S$4.3).
Example 2: the optimization objective of $g_{a}$ is to select an attack rationale $Z_{A}$ from input such that, when $Z_{A}$ is fed into the same predictor $p$, it yields a prediction label flipped from its original label. Continuing the previous example in Figure 2, the generator $g$ selects the “.” from a positive input $X^{1}$ with label $1$ as $Z$. Consequently, the predictor $p$ learns to treat the presence of “.” in $Z$ as an indicative feature for positive classification. On the other hand, the goal of $g_{a}$ is to select an attack rationale $Z_{A}$ from a negative input $X^{0}$ with a label 0 in such a way that, when $Z_{A}$ is fed to the same predictor $p$, the prediction result flips from its original label 0 to $1$. Achieving this objective is straightforward: $g_{a}$ simply needs to mimic $g$ by selecting “.” as $Z_{A}$. This suggests that if $g$ identifies $Z$ from $X^{1}$ as a trivial pattern also present in $X^{0}$, then $g_{a}$ can effortlessly select $Z_{A}=Z$ from $X^{0}$, leading to an easy flip of the prediction label of $Z_{A}$ to $1$ in predictor $p$. On the other hand, if $Z$ is a genuine positive rationale unique to $X^{1}$ and the predictor $p$ classifies it correctly, then $g_{a}$ would be unable to find a positive rationale from the negative input $X^{0}$. Therefore, it is difficult for the predictor $p$ to flip $Z_{A}$’s label from 0 to $1$. Thus, we can leverage the attack generator $g_{a}$ to assist in inspecting and identifying sampling bias. $g_{a}$ may easily find a $Z_{A}$ that flips its predicted label in predictor $p$ from its actual label, indicating the presence of semantically unrelated trivial patterns in $Z$.
@@ A.3 Implementation details of Equation (8) and (9)
Then, we get a copy of $Z_{A}$ with the pytorch function ``torch.detach()":
@@
Note that this updating process will not influence the attacker, since we have used ``torch.detach()" for $Z_{A}$.
@@ A.4 Datasets
The statistics of the datasets are in Table 5. $Pos$ and $Neg$ denote the number of positive and negative examples in each set. $S$ denotes the average percentage of tokens in human-annotated rationales to the whole texts. avg_lenavg\_len denotes the average length of a text sequence.
@@ C.1 Derivation of Equation (13)
The first property is a general property for conditional probability. If $0<P(Y=1)<1$, then for $\forall p$, if $0<p<1$, we can always find a variable $c$, such that $P(Y=1|c)=p$.
@@
If we can construct $G$ in an arbitrary way, and $0<P(Y=1|Z=t)<1$, then we have
@@ C.3 The minimum cross-entropy is equal to entropy
Figure 11: An example of llama’s output. Here “1” means that the class label $Y$ is positive. And the words after “||” represent the rationale.
@@
Appendix E Connection to More Fields
As deep learning progresses, enhancing few-shot learning (Tong et al., 2024b, a, 2025b) and improving the efficiency of multi-modal large language models (MLLMs) (Tong et al., 2025a) are becoming key challenges. We see both as promising directions for extending our work.